convert the following equation to Cartesian coordinates,
r sin 0=-2

Answer by venugopalramana(3286)

convert the following equation to Cartesian coordinates,
r sin 0=-2
HOPE YOU MEAN RSIN(THETA)=-2...LET US USE T FOR THETA...
RSIN(T)=-2
HOPE YOU KNOW THESE 2 EQNS..
RCOS(T)=X
RSIN(T)=Y
HENCE Y=-2 IS THE CARTESIAN FORM
ALITER...IF YOU KNOW THE OTHER FORMULAE...
R = SQRT(X^2+Y^2)
TAN(T)=Y/X
SIN(T)/COS(T)=Y/X
SIN(T)=(Y/X)COS(T)=(Y/X)/SEC(T)=(Y/X)/[SQRT(1+TAN^2(T)]
= (Y/X)/SQRT[1+Y^2/X^2]=(Y/X)/[SQRT(X^2+Y^2)/X]
=Y/SQRT(X^2+Y^2)
HENCE THE GIVEN EQN. IN CARTESIAN COORDINATES IS
SQRT(X^2+Y^2)*Y/SQRT(X^2+Y^2)=Y = -2